OpenStudy (anonymous):

2w + 4x = -7 Been trying to get help for about 2 hours now with no luck, What are the steps to find three solutions to graph for this problem?

OpenStudy (unklerhaukus):

what are you trying to do to that equation?

OpenStudy (unklerhaukus):

are you trying to graph the straight line? to solve for x? to solve for w?

OpenStudy (anonymous):

I'm trying to find three solutions, thats what my textbox tells me to do.

OpenStudy (unklerhaukus):

maybe you should solve for one of the variables first, let's solve for w 2w + 4x = -7 take away 4x from both sides 2w + 4x -4x= -7 -4x 2w = -7 -4x divide both sides by 2 2w/2 = -7/2 -4x/2 w = -7/2 -2x

OpenStudy (unklerhaukus):

now we can make a table